OpenStudy (anonymous):

P=2l+2w. solve for w. is the answer P-l? or is it P/2 -l

OpenStudy (debbieg):

What do you think it is? What steps do you do? P=2L + 2W (easier to read with a capital L) - what is the first step?

OpenStudy (anonymous):

Well I subtracted 2L so I had P-2L then I divided by 2 and that's either P-L or P/2 -L

OpenStudy (debbieg):

Well, if you divide P-2L by 2, you get \(\Large \dfrac{P-2L}{2}\) right? Now you COULD split that up, I supposed, although I don't think it's necessary to do so. But if you do, you get \(\Large \dfrac{P}{2}-\dfrac{2L}{2}=\dfrac{P}{2}-L\) But you definitely do NOT get P-L!! :)